# EEsizer: LLM-Based AI Agent for Sizing of Analog and Mixed Signal Circuit

### v1

@@ V Conclusion
We have developed EEsizer, an LLM-based AI agent that automates the transistor sizing process. The proposed closed-loop methodology enables data collection and annotation throughout optimization process, thereby minimizing manual effort for data collection. Combined with CoT prompt strategies that guide the LLMs step by step, the agent autonomously learns the trade-offs between transistor dimensions and performance, effectively serving as a sizing optimizer without domain-specific fine-tuning or reliance on external algorithms such as BO and gm/Id scripts. These capabilities, which were developed in this work, enable the agent to be adapted to various LLMs, circuits, specifications, and technology nodes without requiring modifications to prompts and functions. As a result, the proposed agent optimized six basic circuits, which include two kinds of amplifiers, a ring oscillator, and three logic gates, and further explored a 20-transistor opamp. Compared to [27], we demonstrated successful node transfer from 180 nm to 90 nm, covering ten performance metrics with rail-to-rail requirements using three LLMs. At 90 nm, OpenAI o3 achieved the user-intended design targets across three test groups within 20 iterations, highlighting its adaptability at advanced nodes for analog circuits. Furthermore, a complete opamp design with a bias circuit was implemented, and 50 variations were performed on transistor size and $V_{th}$, achieving a minimum pass rate of 76% for gain, which demonstrates the robustness of the proposed agent. In conclusion, this agent reflects the potential for optimizing more complex circuits and systems in the future.

### v2

@@ V Conclusion
We have developed EEsizer, an LLM-based AI agent that automates the transistor sizing process. The proposed closed-loop methodology enables data collection and annotation throughout optimization process, thereby minimizing manual effort for data collection. Combined with CoT prompt strategies that guide the LLMs step by step, the agent autonomously learns the trade-offs between transistor dimensions and performance, effectively serving as a sizing optimizer without domain-specific fine-tuning or reliance on external algorithms such as BO and gm/Id scripts. These capabilities, which were developed in this work, enable the agent to be adapted to various LLMs, circuits, specifications, and technology nodes without requiring modifications to prompts and functions. As a result, the proposed agent optimized six basic circuits, which include two kinds of amplifiers, a ring oscillator, and three logic gates, and further explored a 20-transistor opamp. Compared to [27], we demonstrated successful node transfer from 180 nm to 90 nm, covering ten performance metrics with rail-to-rail requirements using three LLMs. At 90 nm, OpenAI o3 achieved the user-intended design targets across three test groups within 20 iterations, highlighting its adaptability at advanced nodes for analog circuits. Furthermore, a complete opamp design with a bias circuit was implemented, and 50 variations were performed on transistor size and $V_{\text{th}}$, achieving a minimum pass rate of 76% for gain, which demonstrates the robustness of the proposed agent. In conclusion, this agent reflects the potential for optimizing more complex circuits and systems in the future.